How much oxygen is dissolved in 100mL water at 298K if partial pressure
of oxygen is 0.5 atm and Kh=1.4*10-3mol/L/atm?

Grade:12

1 Answers

Nida
5 years ago
Concentration = partial pressure ×kH
                      n/V= p ×kH
             n= p×kH× V
 Put values, n= 0.5×1.4×10-3 × 0.1
              n=0.00007
32g of  Oxygen = 0.00007×32
                          = 2.24×10-3
                        =2.24mg